On August 19 you sold the APR 15, K=30, call and at the same time you sold the APR 15 K=30 put. Suppose that both options will not be exercised till their expiration; Calculate your profit/loss per share at expiration if MMML’s price at expiration were:

6.1 S = 35; 6.2    S = 30;      6.3    S = 25.

MMM; TUE August 19 2014. St 27.50 CALLS LAST PUTS LAST Sep14 Oct14Jan15Apr15Sep14Oct14Jan15 Apr15 20 25 27.5 30 32.5 35 37.5 8.50 .35 3.50 3.80 .15 .50 1.35 .55 4.20 1.70 .75 35 .24 2.40 2.75 .45 1.30 5.85 8.10 11.0011.74 8.75 .05 1.00 8.85 9.50 .50 .94 12.50
Answer #1

Cash inflow on Aug 19 = April 15 K=30 call option premium + April 15 K = 30 put option premium i.e. 2.4 + 4 .2 = 6.6

If S = 35

Call option will get exercised in this case and put option premium would be zero

April K = 30 call option outflow at expiration = 35-30 = 5

Therefore profit per share = 6.6 - 5 = 1.6

If S = 30

No option will get exercised in this case.

Therefore profit per share = 6.6 - 0 = 6.6

If S = 25

Put option will get exercised in this case and call option premium would be zero

April K = 30 put option outflow at expiration = 30-25 = 5

Therefore profit per share = 6.6 - 5 = 1.6